OpenStudy (anonymous):

cos,sine,tan??????? On the observation platform in the Statue of Liberty John is approximately 242 feet above ground. He sights a ship in the harbor and measures the angle of depression as 18 degrees.

OpenStudy (anonymous):

If I'm reading this right ...._________________Horizontal distance .....\angle of depression .......\ .........\ ...........\ .............\ ................\ ..................\ ....................\90-42 ......,...............\ Sin42/36 = Sin(90-42)/Horizontal Distance Horizontal Distance(0.67) = (0.74)(36) Horizontal Distance = 39.8 ft.

OpenStudy (anonymous):

he angle of depression is the same as the angle of elevation.) Find the distance from the ship to the statue.

OpenStudy (anonymous):

still use sin? and would be regular sine or sine-1?

OpenStudy (anonymous):

oh follow the above steps!

OpenStudy (anonymous):

i was just asking srry

OpenStudy (anonymous):

No problem, im here 2 help u! You'll get the answer